OpenStudy (anonymous):

s the equation true, false, or open? Explain. 9p + 8 = 10p + 7

OpenStudy (anonymous):

9p + 8 = 10p + 7 First you will subtract 9p from both sides. -9p -9p ----------------- 8 = p + 7 then subtract 7 from both sides. -7 - 7 -------------- 1 = p Its true only when p = 1 Hope this helped!